O Even to Last Response 2. -/0.1 points 0/4 Submissions Used Assume you dissolve 0.235 g of the weak acid benzoic acid, C.HCO please box answers and in understandable handwritng thanks
Answer #1

Males of benzoic acid - may 0.2359 53 = 0.00192 molaumas 122. 12g Amal mot Concentration of benzoic acid = 0.00192X/0 0.0027M0.655x100M x= [on] = [ch₂ Cook] [ut] 10-14 = 1053X10-8 0.655x10-6 Pu= -log[ht] Py = -log (1.53 X108) PH 7.815 Answers a) [130